# Social Bias Probing: Fairness Benchmarking for Language Models

### v2

@@ Abstract
Language models have been shown to encode a variety of social biases, which carries the risk of downstream harms. While the impact of these biases has been recognized, prior methods for bias evaluation have been limited to binary association tests on small datasets, offering a constrained view of the nature of social biases within language models. In this paper, we propose a novel framework for probing language models for social biases. We collect a probing dataset to analyze language models’ general associations, as well as along the axes of social categories, identities, and stereotypes. To this end, we leverage a novel perplexity-based fairness score. We curate a large-scale benchmarking dataset addressing the limitations of existing fairness collections, expanding to a variety of different identities and stereotypes. When comparing our methodology with prior work, we demonstrate that biases within language models are more nuanced than previously acknowledged, suggesting that the scope of biases LMs encode is broader than previously understood. Moreover, we expose how identities expressing different religions lead to the most pronounced disparate treatments across all models.
@@ 1 Introduction
The unparalleled ability of language models to generalize from vast corpora is tinged by an inherent reinforcement of social biases. These biases are not merely encoded within language models’ representations but are also perpetuated to downstream tasks (Blodgett et al., 2021; Stańczak and Augenstein, 2021), where they can manifest in an uneven treatment of different demographic groups – a challenge documented across various studies (Rudinger et al., 2018; Stanovsky et al., 2019; Kiritchenko and Mohammad, 2018; Venkit et al., 2022).
Figure 1: Workflow of Social Bias Probing Framework.
@@
A direct analysis of biases encoded within language models allows to pinpoint the problem at its source, potentially obviating the need for addressing it for every application (Nangia et al., 2020). Therefore, a number of studies have attempted to evaluate social biases within language models (Nangia et al., 2020; Nadeem et al., 2021; Stańczak et al., 2023; Nozza et al., 2022a). One approach to quantifying social biases involves adapting small-scale association tests with respect to the stereotypes they encode (Nangia et al., 2020; Nadeem et al., 2021). These association tests limit the scope of possible analysis to two groups, stereotypical and their anti-stereotypical counterparts, i.e., the identities that “embody” the stereotype and the identities that violate it, often connotated with different/opposite beliefs. This binary approach, which assumes a singular “ground truth” with respect to a stereotypical statement, has restricted the depth of the analysis and simplified the complexity of social identities and their associated stereotypes. The nuanced nature of social biases within language models has thus been largely unexplored.
@@
To overcome these limitations, we introduce a novel probing framework, as outlined in Figure 1. The input of our approach consists of a dataset gathering stereotypes and a set of identities belonging to different social categories: gender, religion, disability, and nationality. First, we combine stereotypes and identities resulting in our probing dataset. Secondly, we assess social biases across five language modeling architectures in English. We propose perplexity (Jelinek et al., 1977), a measure of a language model’s uncertainty, as a proxy for bias. By evaluating how a language model’s perplexity varies when presented with probes that contain identities belonging to different social categories, we can infer which identities are considered the most likely. Using the perplexity-based fairness score, we conduct a three-dimensional analysis: by social category, identity, and stereotype for each of the considered language models. In summary, the contributions of this work are:
@@
  • We conceptually facilitate fairness benchmarking across multiple identities going beyond the binary approach of a stereotypical and an anti-stereotypical identity.
@@
  • We deliver SoFa (Social Fairness), a benchmark resource to conduct fairness probing addressing drawbacks and limitations of existing fairness datasets, expanding to a variety of different identities and stereotypes. [^1] SoFa will be made available upon paper acceptance.
  • We propose a perplexity-based fairness score to measure language models’ associations with various identities.
@@
  • We study social biases encoded within five different language modeling architectures along the axes of social categories, identities, and stereotypes.
@@
A comparative analysis with the popular benchmarks CrowS-Pairs Nangia et al. (2020) and StereoSet Nadeem et al. (2021) reveals marked differences in the overall fairness ranking of the models, suggesting that the scope of biases LMs encode is broader than previously understood. Moreover, we expose how identities expressing religions lead to the most pronounced disparate treatments across all models, while the different nationalities appear to induce the least variation compared to the other examined categories, namely gender and disability.
@@ 2 Related Work
Presenting a recent framing on the fairness of language models, Navigli et al. (2023) define social bias [^2] The term social characterizes bias in relation to the risks and impacts on demographic groups, distinguishing it from other forms of bias, e.g. the statistical one. as the manifestation of “prejudices, stereotypes, and discriminatory attitudes against certain groups of people” through language. Social biases are featured in training datasets and propagated in downstream NLP applications, where it becomes evident when the model exhibits significant errors in classification settings for specific minorities or generates harmful content when prompted with sensitive identities.
Fairness Scores and Datasets
Delobelle et al. (2022) report various bias metrics for pre-trained LMs. The HONEST score, proposed by Nozza et al. (2021), is computed on template-based sentences created to measure the hurtfulness of LMs’ completions within the task of masked language modeling.
Recent work Blodgett et al. (2021) has pointed out relevant concerns regarding stereotype framing and data reliability of benchmark collections explicitly designed to analyze biases in language models, such as CrowS-Pairs Nangia et al. (2020) and StereoSet Nadeem et al. (2021). The scores proposed in the contributions presenting the datasets are highly dependent on the form of the resource they propose and therefore they are hardly generalizable to other datasets to conduct a more general comparative analysis. Specifically, Nangia et al. (2020) leverage pseudo-log likelihood Salazar et al. (2020) based scoring. The score assesses the likelihood of the unaltered tokens based on the modified tokens’ presence. It quantifies the proportion of instances where the LM favors the stereotypical sample (or, vice versa, the anti-stereotypical one). The stereotype score proposed by Nadeem et al. (2021) differs from the former as it allows bias assessment on both masked and autoregressive language models, whereas CrowS-Pairs is limited to the masked ones. Another significant constraint of both datasets is the establishment of a bias score threshold at $50\%$ (Pikuliak et al., 2023). It implies that a model displaying a preference for stereotypical associations more than $50\%$ of the time is considered biased, and vice versa. This threshold implies that a model falling below it may be deemed acceptable or, in other words, unbiased. Further, these datasets exhibit limitations regarding their focus and coverage of identity diversity and the number of stereotypes. This stems from their reliance on the binary comparison between two rigid alternatives — the stereotypical and anti-stereotypical associations — which fails to capture the phenomenon’s complexity. Indeed, they do not account for how the model behaves in the presence of other plausible contexts, and these potential associations need scrutiny for low probability generation by the model, as they can be harmful regardless of the specific target. Therefore, bias measurements using these resources could lead to unrealistic and inaccurate fairness evaluations.
Given the constraints of current solutions, our work introduces a dataset that encompasses a wider range of identities and stereotypes. The contribution relies on a novel framework for probing language models for social biases. To address the limitations identified in the literature review, we design a perplexity-based ranking that produces a more nuanced evaluation of fairness. Regarding adopting of perplexity as a fairness measure, we recognize similarities in the insights leveraged by All Unmasked Likelihood, a novel bias evaluation measure for masked LMs proposed by Kaneko and Bollegala (2022). Compared to AUL and other similar evaluations Nangia et al. (2020); Nadeem et al. (2021), (i) our evaluation approach does not assume but rather rejects the existence of a stereotypical and an anti-stereotypical phrase as framing to measure what is biased and what is not; (ii) we do not differ between modified and unmodified tokens; (iii) we contest the need for a threshold of $50$ considered as a neutral, desirable value; (iv) certain scores are specifically suitable for masked LMs, while we focus our analysis and our metric for autoregressive causal LMs.
@@ 3 Social Bias Probing Framework
The proposed Social Bias Probing framework serves as a technique for fine-grained fairness benchmarking of LMs. Contrary to the existing fairness assessments, which rely on a dichotomous framework of stereotypical vs. anti-stereotypical associations, our methodology expands beyond this oversimplified binary categorization. Ultimately, our approach enables the comprehensive evaluation of language models by incorporating a diverse array of identities, thus providing a more realistic and rigorous audit of fairness within these systems.
@@
Figure 1 presents our workflow. We first collect a set of stereotypes and identities leveraging the Social Bias Inference Corpus (SBIC; Sap et al. 2020) and an identity lexicon curated by Czarnowska et al. (2021). At this stage, we develop the new SoFa (Social Fairness) dataset, which encompasses all probes — identity–stereotype combinations (Section 3.1). The final phase of our workflow involves evaluating language models by employing our proposed perplexity-based fairness measure in response to the constructed probes (Section 3.2).
3.1 Probing Dataset Generation
Our approach requires a set of identities from diverse social and demographic groups, alongside an inventory of stereotypes.
Stereotypes
@@ 3.1 Probing Dataset Generation
We derive stereotypes from the list of implied statements in SBIC, a corpus that collects social media posts having harmful biased implications. The posts are sourced from previously published collections that include English content from Reddit and Twitter, for a total of $44,000$ instances. Additionally, the authors draw from two “Hate Sites”, namely Gab and Stormfront. Annotators were asked to label the texts based on a conceptual framework designed to represent implicit biases and offensiveness. [^3] We refer to the dataset for an in-depth description (https://maartensap.com/social-bias-frames/index.html).
We emphasize that SBIC serves as an exemplary instantiation of our framework. Our methodology can be applied more broadly to any dataset containing stereotypes directed towards specific categories. As not all instances of the original dataset have an annotation regarding the stereotype implied by the social media comment, we filter it to isolate abusive samples having a stereotype annotated. Since certain stereotypes contain the targeted identity, whereas our goal is creating multiple control probes with different identities, we remove the subjects from the stereotypes, performing a preprocessing to standardize the format of statements (details are documented in Appendix A). Finally, we discard stereotypes with high perplexity scores to remove unlikely instances [^4] Other works in the literature also adopt this procedure, e.g., Barikeri et al. (2021).. Since our evaluation is based on identifying and quantifying perplexity peaks (both low and high), keeping statements in the dataset that are unlikely according to the models would offset the results. The filtering is based on a threshold, specifically averaging perplexity scores from each model and creating a histogram to retain only stereotypes in selected bins exhibiting the lowest scores (Fig. 6).
Identities
@@
While we could have directly used the identities provided in the SBIC dataset, we chose not to due to their unsuitability from frequent repetitions and varying expressions influenced by individual annotators’ styles. To unify the set of analyzed identities, we deploy the lexicon created by Czarnowska et al. (2021): Table 5 in Appendix reports a sample for each category. [^5] The complete list of identities is available at https://github.com/amazon-science/generalized-fairness-metrics/tree/main/terms/identity_terms. We map the SBIC dataset target group categories to the identities available in the lexicon (Table 3). Specifically, the categories are: gender, race, culture, disabilities, victim, social, and body. We first define and rename the culture category to include religions and broaden the scope of the race category to encompass nationalities. We then link the categories in the SBIC dataset to those present in the lexicon as follows: gender identities are extracted from the lexicon’s genders and sexual orientations, nationality identities are derived from race and country entries, religion utilizes terms from the religion category, and disabilities are drawn from the disability category. This mapping results in excluding the broader SBIC categories, i.e., victim, social, and body due to the difficulty in aligning the identities with the lexicon categories, and disproving the assumption of invariance in the related statements. [^6] This choice is motivated by the fact that the stereotypes under these categories are often specific to a particular identity; for example, they might have referenced body parts belonging to one gender and not another. The assignment of an identity to a specific category is inherited from the categorizations of the resources adopted. Recognizing that these framings inevitably simplify the complex real-world nuances is crucial.
SoFa
@@
To obtain the final probing dataset, we remove duplicated statements and apply lower-case. Finally, each target is concatenated to each statement with respect to their category, creating dataset instances that differ only for the target (Table 6). In Table 3, we report the coverage statistics regarding targeted categories and identities.
3.2 Fairness Measure
Measure
@@ 3.2 Fairness Measure
We propose perplexity (PPL; Jelinek et al. 1977) as a means of intrinsic evaluation of fairness in language models. PPL is defined as the exponentiated average negative log-likelihood of a sequence. More formally, let $X=(x_{0},x_{1},\dots,x_{t})$ be a tokenized sequence, then the perplexity of the sequence is
@@
where $\log p_{\theta}(x_{d}\mid x_{<d})$ is the log-likelihood of the $d$th token conditioned on the proceeding tokens given a model parametrized with $\theta$.
Our metric leverages PPL to quantify the propensity of a model to produce a given input sentence: a high PPL value suggests that the model deems the input improbable for generation. We identify bias manifestations when a model exhibits low PPL values for statements that contain stereotypes, thus indicating a higher probability of their generation. The purpose of this metric, and more generally our framework, is to provide a fine-grained summary of models’ behaviors from an invariance fairness perspective, i.e., the same statement referring to different demographic groups should not cause a significant change in model behavior, or, in more general terms, individuals from different demographic groups should be treated equally.
[…]
@@
Since the identities $i$ are characterized by their own PPL scores, we normalize the PPL of the probe with the PPL of the identity, addressing the risk that certain identities might yield higher PPL scores because they are considered unlikely.
We highlight that the PPL’s scale across different models can significantly differ based on the training data. Consequently, the raw scores do not allow direct comparisons. We facilitate the comparison of the PPL values of model $m_{1}$ and model $m_{2}$ for a given combination of identity and a stereotype:
@@
In Eq. 2, $k$ is a constant and represents the factor that quantifies the scale of the scores emitted by the model: in fact, different models emit scores having different scales and, therefore, as already mentioned, are not directly comparable. Importantly, each model has its own $k$ [^7] The constant $k$ is not calculated; it is only formally described. The assumption of the existence of this constant $k$ allows us to compare perplexity values through the steps outlined., but because it is a constant, it does not depend on the input text sequence but solely on the model $m$ in question. In Eq. 3, we use the base-$10$ logarithm of the PPL values generated by each model to analyze more tractable numbers since the range of PPL is $[0,\inf)$. For the purpose of calculating variance across the probes $P_{c,s}$ (Eq. 3.2), which is the main investigation conducted in our dataset, $k$ plays no role and does not influence the result. Consequently, we can compare different PPLs from models that have been transformed in this manner.
@@
Lastly, we define Delta Disparity Score (DDS) as the magnitude of the difference between the highest and lowest PPL score as a signal for a model’s bias with respect to a specific stereotype. DDS is computed separately for each stereotype $s$ belonging to category $c$, or, in other words, on the set of probes created from the stereotype $s$.
$DDS_{P_{c,s}}=\max_{P_{c,s}}(\log({PPL^{\star m}_{(i+s)}}))$ (5)
$-\min_{P_{c,s}}(\log({PPL^{\star m}_{(i+s)}}))$
Evaluation
We conduct the following types of evaluation: intra-identities, intra-stereotypes, intra-categories, and calculate a global fairness score. At a fine-grained level, we identify the most associated sensitive identity intra-$i$, i.e., for each stereotype $s$ within each category $c$. This involves associating the $i$ achieving the lowest (top-$1$) $\log({PPL^{\star m}_{(i+s)}})$ as reported in Equation 3, PPL from now on for the sake of brevity. Additionally, we delve into the analysis of stereotypes themselves (intra-$s$), exploring DDS as defined in Equation 5 between the maximum and minimum PPLs obtained for the set of probes generated from $s$ (again, for each $c$, for each $s$ within $c$). This comparison allows us to pinpoint the strongest stereotypes within each category (in the sense of the ones causing the lowest disparity w.r.t. DDS), shedding light on the shared stereotypes across identities. Extending our exploration to the intra-category level, we aggregate and count findings from the intra-identities and stereotypes settings. At a broader level, our goal is to uncover, for each sensitive category, the top-$k$ strongest (low PPL) identities and stereotypes within that category. The findings resulting from the various settings are first investigated separately for each model $m$. In the subsequent analysis, we delve into the overlap among the top-$k$ identities and stereotypes, spanning both within and across model families and scales.

### v3

@@ Abstract
While the impact of social biases in language models has been recognized, prior methods for bias evaluation have been limited to binary association tests on small datasets, limiting our understanding of bias complexities. This paper proposes a novel framework for probing language models for social biases by assessing disparate treatment, which involves treating individuals differently according to their affiliation with a sensitive demographic group. We curate SoFa, a large-scale benchmark designed to address the limitations of existing fairness collections. SoFa expands the analysis beyond the binary comparison of stereotypical versus anti-stereotypical identities to include a diverse range of identities and stereotypes. Comparing our methodology with existing benchmarks, we reveal that biases within language models are more nuanced than acknowledged, indicating a broader scope of encoded biases than previously recognized. Benchmarking LMs on SoFa, we expose how identities expressing different religions lead to the most pronounced disparate treatments across all models. Finally, our findings indicate that real-life adversities faced by various groups such as women and people with disabilities are mirrored in the behavior of these models.
@@ 1 Introduction
The unparalleled ability of language models (LMs) to generalize from vast corpora is tinged by an inherent reinforcement of social biases. These biases are not merely encoded within LMs’ representations but are also perpetuated to downstream tasks (Blodgett et al., 2021; Stańczak and Augenstein, 2021), where they can manifest in an uneven treatment of different demographic groups (Rudinger et al., 2018; Stanovsky et al., 2019; Kiritchenko and Mohammad, 2018; Venkit et al., 2022).
Figure 1: Social Bias Probing framework.
@@
Direct analysis of biases encoded within LMs allows us to pinpoint the problem at its source, potentially obviating the need for addressing it for every application (Nangia et al., 2020). Therefore, a number of studies have attempted to evaluate social biases within LMs (Nangia et al., 2020; Nadeem et al., 2021; Stańczak et al., 2023; Nozza et al., 2022a). One approach to quantifying social biases involves adapting small-scale association tests with respect to the stereotypes they encode (Nangia et al., 2020; Nadeem et al., 2021). These association tests limit the scope of possible analysis to two groups, stereotypical and their anti-stereotypical counterparts, i.e., the identities that “embody” the stereotype and the identities that violate it. This binary approach, which assumes a singular “ground truth” with respect to a stereotypical statement, has restricted the depth of the analysis and simplified the complexity of social identities and their associated stereotypes. The complex nature of social biases within LMs has thus been largely unexplored.
@@
Our Social Bias Probing framework, as outlined in Fig. 1, is specifically designed to enable a nuanced understanding of biases inherent in language models. Accordingly, the input of our approach consists of a set stereotypes and identities. To this end, we generate our probing dataset by combining stereotypes from the Social Bias Inference Corpus (SBIC; Sap et al. 2020) and identities from the lexicon by Czarnowska et al. (2021). In this paper we examine identities belonging to four social categories: gender, religion, disability, and nationality. Secondly, we assess social biases across five state-of-the-art LMs in English. We use perplexity (Jelinek et al., 1977), a measure of language model uncertainty, as a proxy for bias. By analyzing the variation in perplexity when probes feature different identities from diverse social categories, we infer which identities are deemed most likely by a model. This approach facilitates a three-dimensional analysis – by social category, identity, and stereotype—across the evaluated LMs. In summary, the contributions of this work are:
@@
  • We conceptually facilitate fairness benchmarking across multiple identities using our Social Bias Probing framework, going beyond the binary approach of a stereotypical and an anti-stereotypical identity.
@@
  • We introduce SoFa (Social Fairness), a benchmark for fairness probing addressing limitations of existing datasets, including a variety of different identities and stereotypes. [^1] SoFa will be made available upon paper acceptance. See the data statement in App. A.
@@
  • We assess social biases in five autoregressive causal language modeling architectures by examining disparate treatment across social categories, identities, and stereotypes.
@@
A comparative analysis with the popular benchmarks CrowS-Pairs Nangia et al. (2020) and StereoSet Nadeem et al. (2021) reveals marked differences in the overall fairness ranking of the models, providing a different view on the social biases encoded in LMs. We further find how identities expressing religions lead to the most pronounced disparate treatments across all models, while the different nationalities appear to induce the least variation compared to the other examined categories, namely gender and disability. We hypothesize that the increased visibility of religious disparities in language models may stem from recent successful efforts to mitigate racial and gender biases. This underscores the urgency for a comprehensive investigation into biases across multiple dimensions. Additionally, our findings indicate that the LMs reflect the real-life challenges faced by various groups, such as women and people with disabilities.
@@ 2 Related Work
Social Bias Benchmarking
Prior work, such as CrowS-Pairs Nangia et al. (2020) and StereoSet Nadeem et al. (2021), was pioneering in benchmarking models in terms of social biases and harmfulness. However, concerns have been raised regarding stereotype framing and data reliability of benchmark collections designed to analyze biases in LMs Blodgett et al. (2021); Gallegos et al. (2023). Specifically, Nangia et al. (2020) determine the extent to which a masked language model prefers stereotypical or anti-stereotypical responses, while the stereotype score developed by Nadeem et al. (2021) expands this approach to include both masked and autoregressive LMs. A significant limitation of both benchmarks is their use of a 50% bias score threshold, where models are considered biased if they prefer stereotypical associations more than half the time, and unbiased otherwise (Pikuliak et al., 2023). Another approach, which does not rely on choosing one correct answer from two options, is the proposed by Kaneko and Bollegala (2022) All Unmasked Likelihood (AUL) method which predicts all tokens in a sentence, considering multiple correct candidate predictions for a masked token, which is shown to improve accuracy and avoid selection bias. Hosseini et al. (2023) instead leverage pseudo-perplexity Salazar et al. (2020) in combination with a toxicity score to assess the tendency of LMs’ to generate statements distinguished between harmful vs. benevolent.
Our Social Bias Probing framework (i) probes biases across multiple identities without assuming the existence of solely two groups and contests the need for a deterministic threshold for dividing these groups; (ii) is developed with benchmarking social bias in the autoregressive causal LMs in mind.
Social Bias Datasets
Benchmarking social bias is highly reliant on the underlying dataset, i.e., the bias categories, stereotypes, and identities it includes (Blodgett et al., 2021; Delobelle et al., 2022). StereoSet presents over $6$k triplets (for a total of approximately $19$k) crowdsourced instances measuring race, gender, religion, and profession stereotypes, while CrowS-Pairs provides roughly $1.5$k sentence pairs (for a total of $3$k) to evaluate stereotypes of historically disadvantaged social groups. Barikeri et al. (2021) introduce a conversational dataset consisting of $11,873$ sentences generated from Reddit conversations to assess stereotypes between dominant and minoritized groups along the dimensions of gender, race, religion, and queerness.
These datasets cover a limited set of identities and stereotypes. Therefore, bias measurements using these resources could lead to inaccurate fairness evaluations. In fact, Smith et al. (2022b) show that they are able to measure previously undetectable biases with their large-scale dataset of over $450,000$ sentence prompts from two-person conversations. Our SoFa benchmark includes a total of $408$ identities and $11,349$ stereotypes across four social bias dimensions, for a total amount of $1,490,120$ probes, presenting an extensive resource for social bias probing of language models.
@@ 3 Social Bias Probing Framework
Social bias [^2] The term social characterizes bias in relation to the risks and impacts on demographic groups, distinguishing it from other forms of bias, e.g., the statistical one. can be defined as the manifestation through language of “prejudices, stereotypes, and discriminatory attitudes against certain groups of people” (Navigli et al., 2023). These biases are featured in training datasets and are carried over into downstream applications, resulting in, for instance, classification errors concerning specific minorities and the generation of harmful content when models are prompted with sensitive identities (Cui et al., 2024; Gallegos et al., 2023).
@@
To measure the extent to which social bias is present in language models, we propose a Social Bias Probing framework (see Fig. 1) which serves as a technique for fine-grained fairness benchmarking of LMs. We first collect a set of stereotypes and identities (Section 3.1-Section 3.2), which results in the SoFa (Social Fairness) dataset (Section 3.3). The final phase of our workflow involves evaluating language models by employing our proposed perplexity-based fairness measures in response to the constructed probes (Section 3.4), exploited in the designed evaluation setting (Section 3.5).
3.1 Stereotypes
@@
We derive stereotypes from the list of implied statements in SBIC Sap et al. (2020), a corpus of $44,000$ social media posts having harmful biased implications written in English on Reddit and Twitter. Additionally, the authors draw from two widely recognized hate communities, namely Gab [^3] https://gab.com/., a social network popular among nationalists, and Stormfront, [^4] https://www.stormfront.org/forum/. a radical right white supremacist forum. [^5] We refer to the dataset for an in-depth description (https://maartensap.com/social-bias-frames/index.html). We emphasize that SBIC serves as an exemplary instantiation of our framework. Our methodology can be applied more broadly to any dataset containing stereotypes directed towards specific identities.
Professional annotators labeled the original posts as either offensive or biased, ensuring each instance in the dataset contains harmful content. We decide to filter the SBIC dataset to isolate only those abusive samples with explicitly annotated stereotypes. Since certain stereotypes contain the targeted identity, whereas our goal is to create multiple control probes with different identities, we remove the subjects from the stereotypes, to standardize the format of statements. Following prior work (Barikeri et al., 2021), we discard obscure stereotypes with high perplexity scores to remove unlikely instances ensuring accurate evaluation based on perplexity peaks of stereotype–identity pairs. The filtering uses a threshold, averaging perplexity scores across models and removing the highest-scored stereotypes (Fig. 4 in Appendix). We then perform a fluency evaluation of the stereotypes to filter out ungrammatical sentences through the distilbert-base-uncased-CoLA model, [^6] https://huggingface.co/textattack/distilbert-base-uncased-CoLA which determines the linguistic acceptability. Lastly, we remove duplicated stereotypes and apply lower-case. Further details on the preprocessing steps are provided in App. B.
3.2 Identities
@@
While we could have directly used the identities provided in the SBIC dataset, we chose not to due to their unsuitability from frequent repetitions and varying expressions influenced by individual annotators’ styles. To leverage a coherent set of analyzed identities, we deploy the lexicon [^7] The complete list of identities is available at https://github.com/amazon-science/generalized-fairness-metrics/tree/main/terms/identity_terms. created by Czarnowska et al. (2021). In Tab. 3 in the Appendix, we report samples for each category. We map the SBIC dataset group categories to the identities available in the lexicon (Tab. 5 in Appendix). Specifically, the categories from SBIC are gender, race, culture, disabilities, victim, social, and body. We first define and rename the culture category to include religions and broaden the scope of the race category to encompass nationalities. We then link the categories in the SBIC dataset to those present in the lexicon as follows: gender identities are drawn from the lexicon’s genders and sexual orientations, nationality from race and country categories, religion and disabilities directly from their respective categories. This mapping excludes the broader SBIC categories–victim, social, and body–due to alignment challenges with lexicon entries and difficulties in preserving statement invariance. [^8] This choice is motivated by the fact that the stereotypes under these categories are often specific to a particular identity; for example, they might have referenced body parts belonging to one gender and not another. While we inherit the assignment of an identity to a specific category the underlying resources, we recognize that these framings may simplify the complexity of identities.
3.3 SoFa
@@
To obtain SoFa, each target is concatenated to each statement with respect to their category, creating dataset instances that differ only for the target. See Tab. 4 in Appendix for a sample of examples of the generated probes. SoFa consists of a total of 408 coherent identities, over 35k stereotypes, and 1.49mio probes. In Tab. 5 in the Appendix, we report the detailed coverage statistics of SoFa and compare it to existing benchmarks.
To gain an overview of the topics covered by the stereotypes, we conduct a clustering analysis. In Section C.2, we describe the clustering algorithm. Most of the stereotypes are associated with sexualization and violence (over 1000 distinct stereotypes each) with other topics such as family neglect, and racial stereotypes, being mentioned (see Fig. 5 for details). Moreover, we analyze stereotypes under the lens of hate speech analysis, i.e., we quantify how many stereotypes are also instances of hate speech. The majority of stereotypes do not exhibit hate speech features. Indeed, although often the stereotypes do not contain explicitly offensive terms, the underlying intent of the original comment is still harmful, conveying a prejudicial, demeaning perspective. We describe our procedure and results in Section C.3.
3.4 Fairness Measures
@@
We use perplexity (PPL; Jelinek et al. 1977) as a means of intrinsic evaluation of fairness in LMs. PPL is defined as the exponentiated average negative log-likelihood of a sequence. More formally, let $X=(x_{0},x_{1},\dots,x_{t})$ be a tokenized sequence, then the perplexity of the sequence is
@@
where $\log p_{\theta}(x_{d}\mid x_{<d})$ is the log-likelihood of the $d$th token conditioned on the proceeding tokens given a model parametrized with $\theta$. We measure the propensity of a model to produce a given output based on PPL, identifying bias manifestations when a model exhibits low PPL values for statements that contain stereotype-containing statements, suggesting a higher probability of their generation. The purpose of our framework, is to provide a fine-grained summary of models’ behaviors from an invariance fairness perspective, i.e., the same statement referring to different demographic groups should not cause a substantial change in model behavior, or, in more general terms, individuals from different demographic groups should be treated equally.
Formally, let $\mathcal{C}=\{\textit{religion},\textit{gender},\textit{disability},\\ \textit{nationality}\}$ be the set of identity categories; we denote one element of $\mathcal{C}$ as $c$. Further, let $i$ be the identity belonging to a specific category $c$, e.g., Catholics and $s$ be the stereotype belonging to $c$, e.g., are all terrorists. We define $P_{i+s}$ as a singular probe derived by the concatenation of $i$ with $s$, e.g., Catholics are all terrorists, while $P_{c,s}=\{i+s\,|\,i\in c\}$ is the set of probes for $s$ gathering all the controls resulting from the different identities that belong to $c$, e.g., {Catholics are all terrorists; Buddhists are all terrorists; Atheists are all terrorists; …}. Finally, let $m$ be the LM under analysis. The normalized perplexity of a probe is computed as follows:
@@
Since the identities are characterized by their own PPL scores, we normalize the PPL of the probe with the PPL of the identity, addressing the risk that certain identities might yield higher PPL scores because they are considered unlikely.
We highlight that the PPL’s scale across different models can significantly differ based on the training data and, therefore, are not directly comparable. We facilitate the comparison of the PPL values of model $m_{1}$ and model $m_{2}$ for a given combination of identity and a stereotype:
@@
In Eq. 2, $k$ is a constant and represents the factor that quantifies the scale of the scores emitted by the model. Importantly, each model has its own $k$, [^9] The constant $k$ is not calculated; it is only formally described. The assumption of the existence of this constant $k$ allows us to compare perplexity values. but because it is a constant, it does not depend on the input text sequence but solely on the model $m$ in question. In Eq. 3, we use the base-$10$ logarithm of the PPL values generated by each model to analyze more tractable numbers since the range of PPL is $[0,\inf)$. From now on, we call $\log_{10}({PPL^{\star m}_{(i+s)}})$ as $\mathbf{PPL^{\star}}$ for the sake of brevity.
Our proposed perplexity-based SoFa score is based on calculating variance across the probes $P_{c,s}$ (Section 3.4). For this purpose, $k$ plays no role and does not influence the result. Consequently, we can compare the values from different models that have been transformed in this manner.
@@
Lastly, we introduce the Delta Disparity Score (DDS) as the magnitude of the difference between the highest and lowest $PPL^{\star}$ score as a signal for a model’s bias with respect to a specific stereotype. DDS is computed separately for each stereotype $s$ belonging to category $c$, or, in other words, on the set of probes created from the stereotype $s$.
$DDS_{P_{c,s}}=\max_{P_{c,s}}(PPL^{\star})-\min_{P_{c,s}}(PPL^{\star})$ (5)
3.5 Fairness Evaluation
We define and conduct the following four types of evaluation: intra-identities, intra-stereotypes, intra-categories, and calculate a global SoFa score.
Intra-identities ($\mathbf{PPL^{\star}}$)
At a fine-grained level, we identify the most associated sensitive identity intra-$i$, i.e., for each stereotype $s$ within each category $c$. This involves associating the $i$ achieving the lowest (top-$1$) $PPL^{\star}$ as reported in Eq. 3.
Intra-stereotypes (DDS)
We analyze the stereotypes (intra-$s$), exploring DDS as defined in Eq. 5. This comparison allows us to pinpoint the strongest stereotypes within each category, i.e., causing the lowest disparity with respect to the DDS, shedding light on the shared stereotypes across identities.
Intra-categories (SoFa score by category)
For the intra-$\mathbf{c}$ level, to obtain a fairness score for each $m$, for each $c$ and $s$, we compute the variance as formalized in Section 3.4 occurring among the probes of $s$, and average it by the number of $s$ belonging to $c$: $\frac{1}{n}\sum^{n}_{j=1}\sigma^{2}(\log_{10}({PPL^{\star m}_{P_{c,s_{j}}}}))\;\forall s=\{s_{j},\dots,s_{n}\}\in c$. We reference this as SoFa score by category.
Global fairness score (global SoFa score)